OpenStudy (callisto):

Non-interesting integration question #3

OpenStudy (callisto):

Actually... It's the next part of the previous question :S How can we related (d) to the previous part so that we can solve (d) easier?

OpenStudy (amistre64):

looks like we can use the setup in C for something; it looks to have the same kind of elements

OpenStudy (callisto):

I though g(x) = cosx / (1+sinx)^2. ...

OpenStudy (amistre64):

does:\[\frac{cos(pi-x)}{(1+sin(pi-x))^2}=-\frac{cos(x)}{(1+sin(x))^2}\]???

OpenStudy (experimentx):

whoa .. that goes smoothly!!

OpenStudy (callisto):

Which part of the result are you using?

OpenStudy (experimentx):

ln( 1 + e^(cos pi - x)) = ln( 1 + e^-(cos x)) = ln (e^cosx + 1) - ln e^cos x ^^^ the up would be -g(x)

OpenStudy (callisto):

So, it's from (c), in the question...

OpenStudy (experimentx):